OpenStudy (anonymous):

What is the slope m and y-intercept point of the line y = 2x + 1?

OpenStudy (anonymous):

y = 2 x + 1 y = m x + b y = (slope)x + (y intercept) slope = 2 y intercept = 1

OpenStudy (anonymous):

slope m = number in a front of x, so m=2 y-int, means x=0, so y=1